# Gradient Minimization Using Higher Order MethodsYao Ji and Guanghui Lan

### v1

@@ 1.1 Notation and terminology
For convex function $f,$ $\partial f(x)$ denotes set of subdifferential at $x.$ For any real number $s,$ $\lceil s\rceil,\lfloor s\rfloor$ denote the nearest integer to $s$ from above and below. Let $m]\triangleq\{1,\dots,m\}$, with $m\in\mathbb{N}_{+}.$
2 Accelerated Cubic Regularization of the Newton Method
In this section, we consider the unconstrained convex optimization problem (1), where $f$ is a twice differentiable convex function on $\mathbb{R}^{n},$ and it satisfies (2) with $p=2,$ i.e., there exists $L_{3}(f)>0$ such that $\|\nabla^{2}f(x)-\nabla^{2}f(y)\|\leq L_{3}(f)\|x-y\|,$ for all $x,y\in\mathbb{R}^{n}.$
We start with a modified version of the ACNM [23], and show that unlike the ACNM which only outputs function value, the modified ACNM can output a slow gradient norm and the its original fast function value residual simultaneously. Define $T_{M}(x)\coloneqq\arg\min_{y\in\mathbb{R}^{n}}\xi_{2,x}(y)$, where
@@ 2 Accelerated Cubic Regularization of the Newton Method
1: $x_{0}\in\mathbb{R}^{n},\{a_{k}\}_{k\geq 1},$ $a_{k}>0,A_{1}>0.$ Compute $x_{1}=T_{L_{3}}(x_{0})$ and define $f_{1}(x)\coloneqq{f(x_{1})+{\tfrac{1}{\sqrt{L_{3}(f)+M}}\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}}}+\tfrac{C}{6}\|x-x_{0}\|^{3}.$
@@
3:   Compute $\nu_{k}=\operatorname*{arg\,min\,}_{x\in\mathbb{R}^{n}}f_{k}(x).$ Choose $A_{k+1}=A_{k}+a_{k}>0$ and $y_{k}$ $=(1-\alpha_{k})x_{k}+\alpha_{k}\nu_{k},\quad\text{where}\quad\alpha_{k}=\tfrac{a_{k}}{A_{k+1}},\quad a_{k}>0.$
@@
$\mathcal{R}_{k}^{1}:\quad$ $A_{k}f(x_{k})+\textstyle\sum_{j=1}^{k}{A_{j}\tfrac{\left\|\nabla f(x_{j})\right\|^{\frac{3}{2}}}{\sqrt{L+M}}}\leq f_{k}^{*}\equiv\min\limits_{x\in\mathbb{R}^{n}}f_{k}(x).$
$\mathcal{R}_{k}^{2}:\quad$ $f_{k}(x)\leq A_{k}f(x)+\tfrac{(2L+C)\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}}{6}\|x-x_{0}\|^{3}+\tfrac{\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}}{\sqrt{L+M}}.$
Here $\mathcal{R}_{k}^{1}$ shows an lower bound for the minimum function value of the estimating function $f_{k}(x),$ and $\mathcal{R}_{k}^{2}$ establishes an upper bound for $f_{k}(x).$ See Appendix A for the proof. Thus, we have the following convergence guarantee for the modified ACNM.
@@
Let the sequence $\{x_{k}\}^{\infty}_{k=1}$ be generated by Algorithm 1 with the parameters $M=2L_{3}(f),C={12L_{3}(f)}/{(\sqrt{2}-1)^{2}},a_{k}={(k+1)(k+2)}/{2},A_{1}=1,$ then for any $k\geq 1,$ we have
@@
It should be noted that the convergence rate of the gradient norm is slower than that of the function value residual. To achieve a point $\tilde{x}$ such that $\|\nabla f(\tilde{x})\|\leq\varepsilon,$ Algorithm 1 requires $\mathcal{O}(L_{3}(f)^{1/2}\|x_{0}-x^{*}\|/\varepsilon^{1/2})$ iterations, while to achieve the function value residual $f(\tilde{x})-f^{*}\leq\varepsilon,$ Algorithm 1 requires $\mathcal{O}(L_{3}(f)^{1/3}\|x_{0}-x^{*}\|/\varepsilon^{1/3})$ iterations. Note that using a regularization technique could improve the convergence rate to $\mathcal{O}([L_{3}(f)^{1/3}\|x_{0}-x^{*}\|^{2/3}/\varepsilon^{1/3}]\log({L_{3}(f)\|x_{0}-x^{*}\|^{2}}/{\varepsilon}))$ [24]. However, the rates differ by a logarithmic term.
We next outline a third-order AR method that utilizes Algorithm 1 to improve its gradient norm convergence rate to that of the function value residual. Notably, the design of Algorithm 2 is built upon accumulative regularization, where the search points $\{x_{s}\}_{s\in[S]}$ are generated by inexactly solving a third-order proximal mapping defined in (7). Algorithm 2 requires a sequence of regularization parameters $\{\sigma_{s}\}_{s\in[S]}$, and the previous search points $\{x_{s}\}_{s\in[S-1]}$ are accumulated within the regularization term. Unlike the fixed regularization [24], Algorithm 2 uses an accumulative, restart-based scheme: each epoch starts at $x_{s-1}$, and the regularizer adds a new term centered at $x_{s}$ while preserving all earlier terms (rather than centered at $x_{0}$). Such an accumulative regularization is close in spirit to the classic accelerated proximal point method [17].
Observe that the proximal term is cubic. We next develop several properties of this cubic term for use in the subsequent algorithmic analysis. Denote the cubic function as $d_{3}(x)=\|x-x_{0}\|^{3}/3$, for any $x_{0}\in\mathbb{R}^{n}.$ Note that it has a Lipschitz continuous Hessian [32] as follows:
@@
Furthermore, it is a uniformly convex function of degree $3$ with parameter $1/2$ [29].
@@
1: Total number of epochs $S,$ strictly increasing regularization parameters $\{\sigma_{s}\}_{s=0}^{S}$ with $\sigma_{0}=0$ and initial point $x_{0}\in\mathbb{R}^{n}.$
2: $x_{S}.$
3: for $s=1,\dots,S$ do
4:   Set initialization at epoch $s$ to $x_{s-1}.$
5:   Compute an approximate solution $x_{s}$ of the proximal subproblem by running Algorithm 1 with the initialization $x_{s-1}$ for $N_{s}$ iterations. (7) $x_{s}\approx$ $\operatorname*{arg\,min\,}\limits_{x\in\mathbb{R}^{n}}\left\{f_{s}(x)\coloneqq f(x)+\textstyle\sum_{i=1}^{s}\tfrac{(\sigma_{i}-\sigma_{i-1})\|x-{x}_{i-1}\|^{3}}{3}\right\}.$
6: output $x_{S}$
We now present the convergence rate of the third-order AR Algorithm 2.
@@
evaluations of the first-, and second-order information of $f$.
@@
By (7), we have $\|\nabla f(x_{S})\|\leq\|\nabla f_{S}(x_{S})\|+\sum_{i=1}^{S}(\sigma_{i}-\sigma_{i-1})\|x_{S}-{x}_{i-1}\|^{2}.$ We start with bounding $\|\nabla f_{S}(x_{S})\|.$ By the property of the power prox-function in (5), we conclude $f_{S}(\cdot)$ has a Lipschitz continuous Hessian with parameter $L_{3}(f_{S})=L_{3}(f)+4\sigma_{S}.$ For all $s\in[S],$ denote $x_{s}^{*}$ as its exact solution, by Lemma 2.1, we have
(9) $\|\nabla f_{S}(x_{S})\|$ ${\leq}\,{\left[3L_{3}(f_{S})\right]^{\frac{1}{3}}\left(\tfrac{80L_{3}(f_{S})\|x_{S}^{*}-x_{S-1}\|^{3}}{N_{S}^{3}}\right)^{\frac{2}{3}}},$
We proceed with bounding $\|x_{S}^{*}-x_{S-1}\|.$ By optimality conditions of (7) at $x_{s-1}^{*}$ and $x_{s}^{*},$ there holds
@@
where in (a), we used the convexity of $f,$ the uniform convexity of power function in (6), and $\nabla f_{s}(x_{s}^{*})=0,$ in (b) we used Lemma 2.1, in (c), we used the fact $\|x_{s}^{*}-x_{s-1}\|\leq\|x_{s-1}^{*}-x_{s-1}\|,$ and in (d), we substitute the choice of $N_{s}$ in (8). Thus, substituting (2.3) and (8) into (9), we have
@@
Combining it with $\|\nabla f(x_{S})\|\leq\|\nabla f_{S}(x_{S})\|+\sum_{i=1}^{S}(\sigma_{i}-\sigma_{i-1})\|x_{S}-{x}_{i-1}\|^{2},$ we conclude $\|\nabla f(x_{S})\|\leq\varepsilon.$ The total evaluations of the first-, and second-order information of $f$ is bounded by
@@
where in (e), we substitute the choices of $\sigma_{k},N_{k}$ in (8) and used the relation $(a+b)^{1/3}\leq a^{1/3}+b^{1/3},$ for all $a,b\geq 0.$ This concludes the proof.
@@
(i) On the parameter choices: The regularization parameter $\sigma_{s}$ is exponentially increasing, while the epoch length $N_{s}$ is exponentially decreasing, scaling inversely with respect to $\sigma_{s}$, i.e., $N_{s}=\Theta([L_{3}(f)/\sigma_{s}]^{1/3})$. The dominant iterations belongs to the first epoch, which determines the final convergence rate, as the remaining $N_{s}$ are summable. Furthermore, the number of epochs $S$ is chosen such that $4\sigma_{S}$ reaches $L_{3}(f)$, after which the algorithm will terminate.
(ii) On the convergence rate: Observe that the convergence rate in terms of the gradient norm in Proposition 2.2 removes the extra $\log({L_{3}(f)\|x_{0}-x^{*}\|^{2}}/{\varepsilon})$ in [24], and matches the convergence rate of the function value residual for ACNM [23]. Furthermore, the regularization parameter $\sigma_{s}$ increases geometrically as the algorithm approaches the true solution $x^{*}$, which enables the algorithm to use shorter epoch lengths $N_{s}$, thus achieving a faster overall convergence rate.
@@ 3 Gradient Minimization Framework for Structured Convex Problems
In this section, we consider solving the convex composite problem (3). We first introduce a general high-order AR framework, and then in Section 3.1, we study its convergence results when the smooth function part has $\nu$-Hölder continuous $p$-th derivatives. In Section 3.2, we focus on the case when the smooth function part has Lipschitz continuous $p$-th order derivative (i.e., $\nu=1$).
Consider the high-order AR framework as in Algorithm 3.
@@
1: Total number of epochs $S,$ strictly increasing regularization parameters $\{\sigma_{s}\}_{s=0}^{S}$ with $\sigma_{0}=0$ and initial point $x_{0}\in\mathbb{R}^{n}.$
2: $x_{S}.$
3: for $s=1,\dots,S$ do
4:   Compute an approximate solution $x_{s}$ of the proximal subproblem (12) $x_{s}\approx\operatorname*{arg\,min\,}\limits_{x\in\mathbb{R}^{n}}\left\{f_{s}(x)\coloneqq f(x)+\textstyle\sum_{i=1}^{s}\tfrac{\sigma_{i}-\sigma_{i-1}}{p+\nu}\|x-{x}_{i-1}\|^{p+\nu}\right\},$ where $p+\nu\geq 2,$ by running some subroutine $\mathcal{A}$ with the initialization $x_{s-1}$ for $N_{s}$ iterations.
5: output $x_{S}$
Observe that Algorithm 3 shares the acceleration spirit of Algorithm 2, and the power of the regularization term depends on the smoothness level of $g.$ The subroutine $\mathcal{A}$ needs to satisfy some properties for AR to converge, for example, $\mathcal{A}$ outputs a slow gradient norm decay and a fast function value residual decay, which we will specify later in the convergence analysis.
Notice that the regularization terms in the subproblem (12) build upon the power prox-function $d_{p+\nu}:\mathbb{R}^{n}\rightarrow\mathbb{R},$ defined as $d_{p+\nu}(x)\coloneqq 1/{(p+\nu)}\|x-x_{0}\|^{p+\nu},x_{0}\in\mathbb{R}^{n}.$ We review its uniform convexity and $\nu$-Hölder continuity of the $p$-th order derivatives as follows [23, 32]. Notice that we only require $p+\nu\geq 2,$ therefore, AR can cover the first order algorithm for $L$-smooth convex function, i.e., $p=\nu=1,$ and the bounded Hessian case, i.e., $p=2,\nu=0.$
@@
Furthermore, its $p$-th order derivatives satisfies $\nu$- Hölder continuity.
@@
Suppose Algorithm 3 terminates at epoch $S$ with a point $x_{S}^{k},k\geq N_{S},$ then, by the construction of $f_{S},$ and suppose $p+\nu\geq 2,$ we have
@@
We next show that the second term can be properly controlled by the regularization parameters, provided that the inner subroutine satisfies a linear convergence condition $\|x_{s}-x_{s}^{*}\|\leq\|x_{s-1}-x_{s-1}^{*}\|/4,$ which can be satisfied by our subroutines, as we will demonstrate in the next subsections.
Denote $D$ as an upper bound on the distance to the set of optimal solutions, i.e.,
(16) $\text{dist}(x_{0},X^{*})\coloneqq\min_{x^{*}\in X^{*}}\|x_{0}-x^{*}\|\leq D.$
It characterizes the initial optimality condition and will be used frequently in our convergence analysis. We have the following convergence guarantee for (15).
@@
Suppose $p+\nu\geq 2,$ the subroutine can compute $x_{s},$ for all $s\in[S]$, such that $\|x_{s}-x_{s}^{*}\|\leq\|x_{s-1}-x_{s-1}^{*}\|/4,$ and suppose in the last epoch $S$, for all $k\geq N_{S}$, there holds $\|x_{S}^{k}-x_{S}^{*}\|\leq\|x_{S-1}-x_{S-1}^{*}\|/4.$ Then, for all $k\geq N_{S},$ there holds
@@
Similar to the proof of Proposition 2.2, by using the optimality condition of (12), for all $s\geq 1,$ $s\in[S],$ there holds
@@
Furthermore, for the last epoch $S,$ if $p+\nu\geq 2,$ there holds
@@
3.1 Extension from accelerated cubic newton to high-order tensor method
In this subsection, we study AR with tensor methods as subroutines to handle convex functions with $\nu$-Hölder smooth continuous $p$-th derivatives. We make the following assumption regarding the subroutine $\mathcal{A}_{s}\coloneqq\mathcal{A}(f,\{\sigma_{i}\}_{i\leq s},\{x_{i-1}\}_{i\leq s}).$

### v2

@@ 1.1 Notation and terminology
For a convex function $f,$ $\partial f(x)$ denotes the subdifferential at $x$. For any real number $s,$ $\lceil s\rceil$ and $\lfloor s\rfloor$ denote the nearest integers to $s$ from above and below, respectively. Let $m]\triangleq\{1,\dots,m\}$, with $m\in\mathbb{N}_{+}.$
2 Acceleration of the Cubic Regularization of Newton’s Method
In this section, we consider the unconstrained convex optimization problem (1), where $f$ is a twice-differentiable convex function on $\mathbb{R}^{n}$ that satisfies (2) with $p=2,$ i.e., there exists $L_{3}(f)>0$ such that $\|\nabla^{2}f(x)-\nabla^{2}f(y)\|\leq L_{3}(f)\|x-y\|,$ for all $x,y\in\mathbb{R}^{n}.$
We start with a modified version of the ACNM [nesterov2008accelerating], and show that unlike the ACNM, which only outputs the function value, the modified ACNM outputs both a gradient norm bound (with a slower rate) and the original fast function-value residual, simultaneously. Define $T_{M}(x)\coloneqq\arg\min_{y\in\mathbb{R}^{n}}\xi_{2,x}(y)$, where
@@ 2 Acceleration of the Cubic Regularization of Newton’s Method
1: $x_{0}\in\mathbb{R}^{n},\{a_{k}\}_{k\geq 1},$ $a_{k}>0,A_{1}>0.$ Compute $x_{1}=T_{L_{3}(f)}(x_{0})$ and define $f_{1}(x)\coloneqq{f(x_{1})+{\tfrac{1}{\sqrt{L_{3}(f)+M}}\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}}}+\tfrac{C}{6}\|x-x_{0}\|^{3}.$
@@
3:   Compute $\nu_{k}=\operatorname*{arg\,min\,}_{x\in\mathbb{R}^{n}}f_{k}(x).$ Choose $A_{k+1}=A_{k}+a_{k}$ and $y_{k}$ $=(1-\alpha_{k})x_{k}+\alpha_{k}\nu_{k},\quad\text{where}\quad\alpha_{k}=\tfrac{a_{k}}{A_{k+1}},\quad a_{k}>0.$
@@
$\mathcal{R}_{k}^{1}:\quad$ $A_{k}f(x_{k})+\textstyle\sum_{j=1}^{k}{A_{j}\tfrac{\left\|\nabla f(x_{j})\right\|^{\frac{3}{2}}}{\sqrt{L_{3}(f)+M}}}\leq f_{k}^{*}\equiv\min\limits_{x\in\mathbb{R}^{n}}f_{k}(x).$
$\mathcal{R}_{k}^{2}:\quad$ $f_{k}(x)\leq A_{k}f(x)+\tfrac{2L_{3}(f)+C}{6}\|x-x_{0}\|^{3}+\tfrac{\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}}{\sqrt{L_{3}(f)+M}}.$
Here, $\mathcal{R}_{k}^{1}$ shows a lower bound for the minimum function value of the estimating function $f_{k}(x),$ and $\mathcal{R}_{k}^{2}$ establishes an upper bound for $f_{k}(x).$ See Appendix A for the proof. Thus, we have the following convergence guarantee for the modified ACNM.
@@
Let the sequence $\{x_{k}\}^{\infty}_{k=1}$ be generated by Algorithm 1 with the parameters $M=2L_{3}(f),C={12L_{3}(f)}/{(\sqrt{2}-1)^{2}},a_{k}={(k+1)(k+2)}/{2},A_{1}=1,$ then, for any $k\geq 1,$ we have
@@
Note that the convergence rate of the gradient norm is slower than that of the function-value residual. To achieve a point $\tilde{x}$ such that $\|\nabla f(\tilde{x})\|\leq\varepsilon,$ Algorithm 1 requires $\mathcal{O}(L_{3}(f)^{1/2}\|x_{0}-x^{*}\|/\varepsilon^{1/2})$ iterations, while to achieve the function-value residual $f(\tilde{x})-f^{*}\leq\varepsilon,$ Algorithm 1 requires $\mathcal{O}(L_{3}(f)^{1/3}\|x_{0}-x^{*}\|/\varepsilon^{1/3})$ iterations. Note that using a regularization technique could improve the convergence rate to $\mathcal{O}([L_{3}(f)^{1/3}\|x_{0}-x^{*}\|^{2/3}/\varepsilon^{1/3}]\log({L_{3}(f)\|x_{0}-x^{*}\|^{2}}/{\varepsilon}))$ [nesterov2012make]. However, the rates differ by a logarithmic term.
We next outline a third-order AR method that utilizes Algorithm 1 to improve its gradient norm convergence rate to that of the function-value residual. Notably, the design of Algorithm 2 is built upon accumulative regularization, where the search points $\{x_{s}\}_{s\in[S]}$ are generated by inexactly solving a third-order proximal mapping defined in (7). Algorithm 2 requires a sequence of regularization parameters $\{\sigma_{s}\}_{s\in[S]}$, and the previous search points $\{x_{s}\}_{s\in[S-1]}$ are accumulated within the regularization term. Unlike the fixed regularization [nesterov2012make], Algorithm 2 uses an accumulative, restart-based scheme: each epoch starts at $x_{s-1}$, and the regularizer adds a new term centered at $x_{s}$ while preserving all earlier terms (rather than centered at $x_{0}$). Such accumulative regularization is close in spirit to the classic accelerated proximal point method [guler1992new].
Observe that the proximal term is cubic. We next develop several properties of this cubic term for use in the subsequent algorithmic analysis. Denote the cubic function as $d_{3}(x)=\|x-x_{0}\|^{3}/3$, for any $x_{0}\in\mathbb{R}^{n}.$ Note that it has a Lipschitz continuous Hessian [rodomanov2020smoothness] as follows:
@@
Furthermore, it is uniformly convex of degree $3$ with parameter $1/2$ [nesterov2018lectures].
@@
1: Total number of epochs $S;$ strictly increasing regularization parameters $\{\sigma_{s}\}_{s=0}^{S}$ with $\sigma_{0}=0;$ initial point $x_{0}\in\mathbb{R}^{n}.$
2: for $s=1,\dots,S$ do
3:   Set initialization at epoch $s$ to $x_{s-1}.$
4:   Compute an approximate solution $x_{s}$ of the proximal subproblem by running Algorithm 1 with the initialization $x_{s-1}$ for $N_{s}$ iterations. (7) $x_{s}\approx$ $\operatorname*{arg\,min\,}\limits_{x\in\mathbb{R}^{n}}\left\{f_{s}(x)\coloneqq f(x)+\textstyle\sum_{i=1}^{s}\tfrac{(\sigma_{i}-\sigma_{i-1})\|x-{x}_{i-1}\|^{3}}{3}\right\}.$
5: return $x_{S}$
We now present the convergence rate of the third-order AR method Algorithm 2.
@@
evaluations of the first- and second-order information of $f$.
@@
By (7), we have $\|\nabla f(x_{S})\|\leq\|\nabla f_{S}(x_{S})\|+\sum_{i=1}^{S}(\sigma_{i}-\sigma_{i-1})\|x_{S}-{x}_{i-1}\|^{2}.$ We start with bounding $\|\nabla f_{S}(x_{S})\|.$ By the property of the power prox-function in (5), we conclude that $f_{S}(\cdot)$ has a Lipschitz continuous Hessian with parameter $L_{3}(f_{S})=L_{3}(f)+4\sigma_{S}.$ For all $s\in[S],$ let $x_{s}^{*}$ denote the exact solution. By Lemma 2.1,
(9) $\|\nabla f_{S}(x_{S})\|$ ${\leq}\,{\left[3L_{3}(f_{S})\right]^{\frac{1}{3}}\left(\tfrac{80L_{3}(f_{S})\|x_{S}^{*}-x_{S-1}\|^{3}}{N_{S}^{3}}\right)^{\frac{2}{3}}}.$
We proceed with bounding $\|x_{S}^{*}-x_{S-1}\|.$ By optimality conditions of (7) at $x_{s-1}^{*}$ and $x_{s}^{*},$ the following hold
@@
where in (a), we used the convexity of $f,$ the uniform convexity of the power prox-function in (6), and $\nabla f_{s}(x_{s}^{*})=0,$ in (b) we used Lemma 2.1, in (c), we used the fact $\|x_{s}^{*}-x_{s-1}\|\leq\|x_{s-1}^{*}-x_{s-1}\|,$ and in (d), we substituted the choice of $N_{s}$ in (8). Thus, substituting (2.3) and (8) into (9), we have
@@
Combining this with $\|\nabla f(x_{S})\|\leq\|\nabla f_{S}(x_{S})\|+\sum_{i=1}^{S}(\sigma_{i}-\sigma_{i-1})\|x_{S}-{x}_{i-1}\|^{2},$ we conclude $\|\nabla f(x_{S})\|\leq\varepsilon.$ The total number of first- and second-order evaluations of $f$ is bounded by
@@
where in (e), we substituted the choices of $\sigma_{k},N_{k}$ in (8) and used the relation $(a+b)^{1/3}\leq a^{1/3}+b^{1/3},$ for all $a,b\geq 0.$ This concludes the proof.
@@
(i) On the parameter choices: The regularization parameter $\sigma_{s}$ is exponentially increasing, while the epoch length $N_{s}$ is exponentially decreasing, scaling inversely with respect to $\sigma_{s}$, i.e., $N_{s}=\Theta([L_{3}(f)/\sigma_{s}]^{1/3})$. The dominant iterations belong to the first epoch, which determines the final convergence rate, as the remaining $N_{s}$ are summable. Furthermore, the number of epochs $S$ is chosen such that $4\sigma_{S}$ reaches $L_{3}(f)$, after which the algorithm terminates.
(ii) On the convergence rate: Observe that the convergence rate in terms of the gradient norm in Proposition 2.2 removes the extra $\log({L_{3}(f)\|x_{0}-x^{*}\|^{2}}/{\varepsilon})$ in [nesterov2012make], and matches the convergence rate of the function-value residual for ACNM [nesterov2008accelerating]. Furthermore, the regularization parameter $\sigma_{s}$ increases geometrically as the algorithm approaches the true solution $x^{*}$, which enables the algorithm to use shorter epoch lengths $N_{s}$, thus achieving a faster overall convergence rate.
@@ 3 Gradient Minimization Framework for Structured Convex Problems
In this section, we study the convex composite problem (3). We first introduce a general high-order AR framework, and then in Section 3.1 we study its convergence when the smooth part has $p$-th derivatives that are $\nu$-Hölder continuous. In Section 3.2, we focus on the case when the smooth function part has Lipschitz continuous $p$-th-order derivative (i.e., $\nu=1$).
Consider the high-order AR framework in Algorithm 3.
@@
1: Total number of epochs $S;$ strictly increasing regularization parameters $\{\sigma_{s}\}_{s=0}^{S}$ with $\sigma_{0}=0;$ initial point $x_{0}\in\mathbb{R}^{n}.$
2: for $s=1,\dots,S$ do
3:   Compute an approximate solution $x_{s}$ of the proximal subproblem (12) $x_{s}\approx\operatorname*{arg\,min\,}\limits_{x\in\mathbb{R}^{n}}\left\{f_{s}(x)\coloneqq f(x)+\textstyle\sum_{i=1}^{s}\tfrac{\sigma_{i}-\sigma_{i-1}}{p+\nu}\|x-{x}_{i-1}\|^{p+\nu}\right\},$ where $p+\nu\geq 2,$ by running some subroutine $\mathcal{A}$ with the initialization $x_{s-1}$ for $N_{s}$ iterations.
4: return $x_{S}$
Observe that Algorithm 3 shares the acceleration idea of Algorithm 2, and the power of the regularization term depends on the smoothness level of $g.$ The subroutine $\mathcal{A}$ needs to satisfy some properties for AR to converge, for example, $\mathcal{A}$ outputs a slow gradient norm decay and a fast function-value residual decay, which we will specify later in the convergence analysis.
Notice that the regularization terms in the subproblem (12) build upon the power prox-function $d_{p+\nu}:\mathbb{R}^{n}\rightarrow\mathbb{R},$ defined as $d_{p+\nu}(x)\coloneqq 1/{(p+\nu)}\|x-x_{0}\|^{p+\nu},x_{0}\in\mathbb{R}^{n}.$ We recall its uniform convexity and the $\nu$-Hölder continuity of the $p$-th-order derivatives as follows [nesterov2008accelerating, rodomanov2020smoothness]. Note that we only require $p+\nu\geq 2.$ Therefore, AR covers the first order algorithm for $L$-smooth convex function, i.e., $p=\nu=1,$ and the bounded Hessian case, i.e., $p=2,\nu=0.$
@@
Furthermore, its $p$-th-order derivatives satisfies $\nu$- Hölder continuity.
@@
Suppose Algorithm 3 terminates at epoch $S$ with a point $x_{S}^{k},k\geq N_{S}.$ Then, by the construction of $f_{S},$ and suppose $p+\nu\geq 2,$ we have
@@
We next show that the second term is controlled by the regularization parameters, provided the inner subroutine satisfies the linear convergence condition $\|x_{s}-x_{s}^{*}\|\leq\|x_{s-1}-x_{s-1}^{*}\|/4,$ which can be satisfied by our subroutines, as we will demonstrate in the next subsections.
Let $D$ denote an upper bound on the distance to the set of optimal solutions, i.e.,
(16) $\operatorname{dist}(x_{0},X^{*})\coloneqq\min_{x^{*}\in X^{*}}\|x_{0}-x^{*}\|\leq D.$
This quantity characterizes the initial optimality gap and will be used frequently in our convergence analysis. We have the following convergence guarantee for (15).
@@
Suppose $p+\nu\geq 2,$ the subroutine can compute $x_{s},$ for all $s\in[S]$, such that $\|x_{s}-x_{s}^{*}\|\leq\|x_{s-1}-x_{s-1}^{*}\|/4,$ and suppose in the last epoch $S$, for all $k\geq N_{S}$, we have $\|x_{S}^{k}-x_{S}^{*}\|\leq\|x_{S-1}-x_{S-1}^{*}\|/4.$ Then, for all $k\geq N_{S},$ we have
@@
Similar to the proof of Proposition 2.2, by using the optimality condition of (12), for all $s\geq 1,$ $s\in[S],$ we have
@@
Furthermore, for the last epoch $S,$ if $p+\nu\geq 2,$ we have
@@
3.1 Extension from accelerated cubic Newton to high-order tensor methods
In this subsection, we study AR with tensor methods as subroutines to handle convex functions with $p$-th derivatives that are $\nu$-Hölder continuous. We make the following assumption regarding the subroutine $\mathcal{A}_{s}\coloneqq\mathcal{A}(f,\{\sigma_{i}\}_{i\leq s},\{x_{i-1}\}_{i\leq s}).$